OpenStudy (anonymous):

Find four consecutive integers whose sum is -38 A. -13, -12, -11, -10 B. -12, -11, -10, -9 C. -11, -10, -9, -8 D. -10, -9, -8, -7

OpenStudy (anonymous):

wow...... Why did i even ask this question?!?!? Im dumb lol

OpenStudy (anonymous):

I know the answer

OpenStudy (anonymous):

the integers are x x+1 x+2 x+3 so x + x+1 + x+2 + x+3 = -38 4x+6 = -38 4x = -44 x = -11 x+1 = -10 x+2 =-9 x+3 = -8